Suppose you purchase a 30-year zero-coupon bond with a yield to maturity of 5.8%. You hold the bond for five years before selOPTION A, B, C AND D THANKS!
Answer #1

a. Price of Bond today =1000/(1+5.8%)^30 =184.2595
After 5 years bond price =1000/(1+5.8%)^25 =244.2633
Rate of Return over 5 years =(244.2633/184.2595)-1=32.56%
Rate of Return of investment per year =(244.2633/184.2595)^(1/5)-1 =5.80%

b. Price of Bond today =1000/(1+6.8%)^30 =138.9512
After 5 years bond price =1000/(1+6.8%)^25 =193.0717
Rate of Return over 5 years =(193.0717/138.9512)-1=38.95%
Rate of Return of investment per year =(193.0717/138.9512)^(1/5)-1 =6.80%

c.  Price of Bond today =1000/(1+4.8%)^30 =244.9974
After 5 years bond price =1000/(1+4.8%)^25 =309.7191
Rate of Return over 5 years =(309.7191/244.9974)-1=26.42%
Rate of Return of investment per year =(309.7191/244.9974)^(1/5)-1 =4.80%

d. No reinvestment risk might not be zero because if bond is redeemed then the return of next investment might not provide the same return . If return is low then the risk increases.